Dublin University Rifle Club 50m Prone Rifle Open 2001 Dublin University Rifle Club 50m Prone Rifle Open 2001 Dublin University Rifle Club 50m Prone Rifle Open 2001

Dublin University Rifle Club 50m Prone Rifle Open 2001 Dublin University Rifle Club 50m Prone Rifle Open 2001 Dublin University Rifle Club 50m Prone Rifle Open 2001

Dublin University Rifle Club 50m Prone Rifle Open 2001 Dublin University Rifle Club 50m Prone Rifle Open 2001 Dublin University Rifle Club 50m Prone Rifle Open 2001

Dublin University Rifle Club 50m Prone Rifle Open 2001 Dublin University Rifle Club 50m Prone Rifle Open 2001 Dublin University Rifle Club 50m Prone Rifle Open 2001

Dublin University Rifle Club 50m Prone Rifle Open 2001 Dublin University Rifle Club 50m Prone Rifle Open 2001 Dublin University Rifle Club 50m Prone Rifle Open 2001

Dublin University Rifle Club 50m Prone Rifle Open 2001 Dublin University Rifle Club 50m Prone Rifle Open 2001 Dublin University Rifle Club 50m Prone Rifle Open 2001

Dublin University Rifle Club 50m Prone Rifle Open 2001 Dublin University Rifle Club 50m Prone Rifle Open 2001 Dublin University Rifle Club 50m Prone Rifle Open 2001

Dublin University Rifle Club 50m Prone Rifle Open 2001 Dublin University Rifle Club 50m Prone Rifle Open 2001 Dublin University Rifle Club 50m Prone Rifle Open 2001

Dublin University Rifle Club 50m Prone Rifle Open 2001 Dublin University Rifle Club 50m Prone Rifle Open 2001 Dublin University Rifle Club 50m Prone Rifle Open 2001